A cob Douglas production function for a firm is given as Q=4L ½K½. The firm has also established that wage rate and interest paid on capital are $3 and $5 respectively for a production period. The firm intents to spend $200 million for the period on production cost. Compute the levels of capital and labor that will maximize output. What is the maximum output?



1
Expert's answer
2021-03-15T20:48:25-0400

Form a Lagragian equation

"Q=4L^{0.5}K^{0.5}" Subject to "wL+rK=C"

"L=4L^{0.5}K^{0.5}-\u03bb(wL+rK-C)"


"\frac{\delta{L}}{\delta{L}}=2L^{-0.5}K^{0.5}-\lambda{w}=0........(i)"


"\frac{\delta{L}}{\delta{K}}=2L^{0.5}K^{-0.5}-\lambda{r}=0........(ii)"


"\frac{\delta{L}}{\delta{\lambda}}=wL+rK-C=0........(iii)"


Divide equation (i) and (ii)


"\frac{K}{L}=\frac{w}{r}" and thus "K=\frac{wL}{r}" and "L=\frac{rK}{w}"


Replacing the two equation on equation (iii)


"w(\frac{kr}{w})+rK=C" thus "K^*=\frac{C}{2r}"


"K^*= \frac{200}{2}×5 =20"


"wL+r(\frac{wL}{r})=C"


"L^*=\frac{C}{2w}"


"L^*=\frac{200}{2}×3 = 33.33"

Hence the optimum output is;

"Q=4(33.33)^{0.5}(20)^{0.5}"


"Q=4(4.58)(4.5)"


Q=104.40 units - Maximum Output.